OpenStudy (payton13579):

The radius of a circle is 17 feet. Using the formula A = r2, what is the area of the circle? Use 3.14 for pi.

Vocaloid (vocaloid):

replace "r" with 17. replace "pi" with 3.14. then calculate A = pi * r^2

OpenStudy (mathmale):

Payton: It's important that you write the equation for the aera of a circle of radius r as A= pi r^2 or \[A=\pi r^2\]

OpenStudy (mathmale):

r2 just doesn't cut it in math. Your formula does not include pi. The correct formula for the area of a circle of radius r is \[A= \pi r^2\] From here, follow Vocoloid's advice. Show your work, please.

OpenStudy (payton13579):

ok

OpenStudy (payton13579):

a=3.14*17 ..

OpenStudy (payton13579):

i forgot how to square it

OpenStudy (payton13579):

@bmk614

OpenStudy (payton13579):

@Vocaloid

OpenStudy (payton13579):

3.14*17=53.38

Vocaloid (vocaloid):

you just multiply that by 17 again

OpenStudy (payton13579):

kk

OpenStudy (payton13579):

907.46

OpenStudy (mathmale):

The square of 17 is simply 17*17 = 289. Use 3.14 as the value of pi. Show your work. What is the area of a circle of radius 17 feet? It's important that you include the units of measurement with your numerical answer.